OpenStudy (anonymous):

help please urgent Integral of 2sin(3x) + e^3x - 1dx

OpenStudy (johnweldon1993):

By parts? Just separate it. \[\large \int 2sin(3x) + e^{3x} - 1 dx\] becomes \[\large \int 2sin(3x)dx + \int e^{3x}dx - \int 1dx\]

OpenStudy (johnweldon1993):

Remember all constants can be moved outside the integral so we have \[\large 2\int sin(3x)dx + \int e^{3x}dx - 1\int dx\]

OpenStudy (johnweldon1993):

And it looks like some simple u-subs will do the trick :)

OpenStudy (anonymous):

\[2-\cos(3x)\div3+e^3x/3-1x+c is this correct?\]

OpenStudy (johnweldon1993):

\[\large -\frac{2}{3}cos(3x) + \frac{1}{3}e^{3x} - x + c\] indeed!

OpenStudy (anonymous):

excellent thanks so very much xoxo

OpenStudy (johnweldon1993):

Anytime :)